Amana owns two houses in the same community. The ratio of the values of the first to the second house is 10: 15 respectively. It is estimated that in 6 years, the value of the first house will increase by 35% and that of the second house will increase by $30,825.00. If the new ratio is 3: 4, find the original value of the first house.
Answer by mananth(16946)

Amana owns two houses in the same community. The ratio of the values of the first to the second house is 10: 15 respectively. It is estimated that in 6 years, the value of the first house will increase by 35% and that of the second house will increase by $30,825.00. If the new ratio is 3: 4, find the original value of the first house.
The ratio of the values of the first to the second house is 10: 15 respectively.
Let the common multiple be x
Values of the first house will be 10x
the value of second house will be 15x
Value of the first house after 6 years will be 1.35* 10x
Value of the second house after 6 years will be 15x +30825.00
Now the ratio is 3:4
(1.35* 10x)/(15x +30825.00) =3/4
13.5x /(15x +30825.00)= 3/4
4* 13.5x = 3*(15x +30825.00)
54x = 45x +92475
9x = 92475
x =10275
Value of first house = 10x = 102,750
Value of second house = 15x = 154125
CHECK
(102750+ (1.35*10275))/(154125+30) =0.756


The original value of the first house is $ 102,750.00